OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $400. How many sophomores bought tickets?

OpenStudy (anonymous):

s = sophomores f = freshmen f + s = 150 2f + 3s = 400 f = 150 - s 2(150 - s) + 3s = 400 300 - 2s + 3s = 400 -2s + 3s = 100 s = 100 100 Sophomores bought tickets.